# Physical properties of circumnuclear ionising clusters. III. Kinematics of gas and stars in NGC 7742

### v2

@@ 3.3 Stellar absorption lines
The main objective of our work is to estimate the properties of the stellar population of the clusters associated to the observed CNSFR. In order to do that we have used two sets of absorption lines: the MgIb triplet at 5167-5183 detected in the MUSE data (unresolved) and the CaII triplet (CaT) at 8498, 8542, 8662 Å (CaT) resolved with MEGARA.
We have measured the equivalent widths (EW) of these stellar absorption lines by integrating the line fluxes and taking as continuum values the linear interpolation between the mean integrated fluxes in 30 Å continuum bands at both sides of the line. The details of the measuring procedure can be seen in Tab. 8 of Zamora & Díaz (2023a). The EW errors have been calculated using the standard deviation in the two continuum bands and propagating them in quadrature. In principle, absorption line EW measurements in objects with different velocity dispersions should be for the broadening of the spectral lines; this effect decreases the continuum level providing lower EW values. However, this correction for MUSE data was found to be negligible (¡ 1 Å for velocity dispersions of around 170 km/s for MUSE spectral resolution), then no correction has been applied. Table 5 lists in columns 1 to 3: (1) the region ID; (2) the CaT EW in Å; and (3) the MgIb EW also in Å .
@@
a (8542Å + 8662Å+ 8498Å).
The presence of a non negligible contribution by the nebular continuum can dilute the starlight weakening and distorting the absorption features. This component is usually not taken into account, but it provides information about the origin of the features measured. Thus, we have calculated a dilution factor, D, for each set of lines with respect to a reference average value measured in the spectra of normal spiral galaxy nuclei which corresponds to old, metal rich stellar populations (EWref(CaII) = 7.7 0.5 Å, EWref(MgI) = 5.18 0.71 Å; Terlevich et al., 1990; Keel, 1983). We have calculated this dilution factor as the ratio between the measured EWs and the adopted reference values: D = EWobs / EWref and compared it to that dilution produced by the nebular continuum associated with a nebula ionised by a young star cluster synthesised using the PopStar code (Mollá et al., 2009) with Salpeter’s IMF (mlow = 0.85 M⊙, mup = 120 M⊙ Salpeter, 1955) and an age of 5.5 Ma.
Figure 4 shows the comparison between the dilution factors, D, calculated for the MgIb and CaT stellar absorption features for the studied CNSFRs. The red lines represent the dilution produced by the nebular continuum in the optical and near IR. We can see that the mean value of the MgIb dilution is about 50%, consistent with the contribution by a nebular continuum. However, the CaT lines are shown almost undiluted. This result can be explained by the presence of red supergiant stars whose CaT features are stronger than those in normal early spiral galaxies while the reference value probably consists of a stellar population dominated by red giants. This points to the occurrence of a relatively recent star formation episode. The same behaviour has already been found for circumnuclear star clusters in both rings of NGC 7469 (see Zamora & Díaz, 2023a) and in earlier work (Terlevich et al., 1990; Garcia Vargas et al., 1993; Oliva et al., 1995).
@@ 3.4 Stellar velocity dispersions
Stellar velocity dispersions can be obtained from the stellar absorption lines. In the second paper of this series (see Zamora & Díaz, 2023a) we used the CaT lines detected in MUSE data for this calculation. However, the spectral resolution reached in MUSE data was insufficient to resolve the CaT lines in the case of the HII ring regions in NGC 7742 and therefore we acquired higher resolution data using MEGARA which provides about 14 km/s of spectral resolution (see Sec. 2). As in the case of NGC 7469, the clusters studied in this work have red supergiant stars dominating the near IR range of the spectrum and the CaT 8498, 8542, 8662 Å lines (see Sec. 3.3) are easily detected. The use of the near IR spectra provides two principal advantages for this study: the contamination by TiO bands and nebular lines is small and the velocity resolution at longer wavelengths is higher than in the blue part of the spectrum for the same spectral dispersion. Additionally, using the cross correlation technique, stellar weak lines present in this part of the spectra also contribute to the goodness of the results.
@@
We have used the cross-correlation technique proposed by Tonry & Davis (1979). In this traditional method, a comparison between a stellar template with the observed stellar population spectrum is done. The assumption made assumes that a galaxy spectrum can be represented by the sum of different stellar spectra with different velocity offsets convolved with a broadening function: , where b is the broadening function, g and t are the cluster and the stellar template spectra respectively, and is the velocity shift between stars. The width of this broadening function is calculated using two correlations: the cluster spectrum with the stellar template (g t) and the stellar template with itself (t t). The convolution products are applied assuming a periodic spectrum with discrete Fourier transforms and the broadening function is assumed to be Gaussian. However, this last assumption is not correct for the (t t) correlation at spectral resolutions as high as the one provided by MEGARA and this technique needs to be modified. We have used the adapted methods proposed in Zamora & Díaz (2023c) of the (g g) correlation which is broader than the other two ones and whose Gaussian behavior can be ensured. In order to test these two additional methods and the traditional ones with our data, we have selected a region with high S/N (R52, located in the pointing with the higher exposure time in the HR-I setup). For the stellar template, we have used late-type red giants and supergiants obtained from the MEGARA stellar library (García-Vargas et al., 2020) aligning all available stellar spectra in velocity and calculating the average between them, verifying that no apparent broadening is introduced in the procedure. The use of only two stellar types of stars for the correlation analysis could introduce errors in our velocity dispersion measures. However, the use of CaT lines minimises the mismatch between the stellar template and the cluster spectrum since they are very strong in most stars of moderate to high metallicity.
@@
The upper panel of Figure 5 shows the stellar spectrum of the selected cluster R52 at high and low MEGARA spectral resolutions, with red and purple lines respectively. The stellar template is also shown with a blue line and the nebular lines present in this spectral range are marked. We can see that the S/N of these spectra is low, even in this region which was observed with twice the average exposure time. This could be due to the noise component or to errors associated with the subtraction of the sky background which is very prominent in the near IR. Thus, we have calculated a mean spectrum of the sky, using the MEGARA fibers earmarked for this purpose. The lower panel of the figure shows this spectrum that shares some features with the cluster one.
@@
We have used the modified version of Tonry & Davis’ method described above for the application of the cross-correlation technique to the data obtained from the MEGARA in the HR-I configuration in the spectral range 8450 Åto 8850 Å. Prior to this, we have binned the spectrum into 4069 logarithmic wavelength bins corresponding to a velocity resolution of 3.4 km/s. Then, we have estimated the continuum by fitting a second order polynomial after masking the features present in the spectrum. The masks have been built assuming a width of 3 Å and 10 Å at each side of the central wavelength of the nebular and stellar lines involved respectively. This continuum has subsequently been subtracted.
@@
Next, the high and low frequency variations have been filtered using a band-pass filter in the Fourier spectrum transform. The minimum and maximum wave numbers are kmin = 3 and kmax = 350 which correspond to wavelength values lower than 10 Å and to the nominal MEGARA/HR-I spectral resolution. The low frequency variations are associated with continuum subtracted errors and the high ones with the noise component. Figure 6 shows, in a small range of wavelengths, in light red the original spectrum and in dark red the spectrum after the band-pass filter. The stellar template is also shown in blue for comparison. We can see that this filter removes part of the noise, recovering some stellar features. However, the S/N continues to be low and we can see residuals for example in the central part of the strongest line. Thus, the noise present in our spectrum probably comes from the errors associated with the pipeline sky subtraction. Therefore, we have decided to use the data without the sky subtraction to be able to quantify this additional component in the cross correlation technique and obtain a more reliable result in the stellar dispersion determination.
@@
An important step in the cross correlation procedure is the subtraction of the non-stellar component of the spectra, which can introduce errors in the velocity dispersions of up to 5 km/s (see Zamora & Díaz, 2023c). In order to remove correctly the nebular component of the spectrum, we have first removed the proper motion of the cluster by shifting the spectrum to the rest frame position. For this step, we have used the position of the (g t) cross correlation. Then, we have masked the nebular lines shown in Fig. 5 assuming a width of 0.4 Å at each side of their central wavelength.
@@
Finally, we have calculated four cross correlation functions for this HII region: (t t), (g t), (g g) and (s s), with t being the stellar template, g the cluster spectrum and s the mean sky spectrum of our observations. The upper left panel of Figure 7 shows the (s s) correlation. Two components can be seen, one associated with the sky lines and another associated with the detector response. They have been fitted with two Gaussian functions and their widths are = 0.2098 Å and = 12.1445 Å respectively. The width of the first component corresponds to a value = 0.15 Å in -units, which is slightly lower than the MEGARA resolution for this setup, = 0.18.
@@
The upper right panel of Figure 7 shows, as an example, the (g g) correlation for the R52 cluster. In this figure we can identify the component associated with the sky lines and the component associated with the stars of the cluster. According to this, we have fitted simultaneously two Gaussian functions, one of them using the value calculated for the (s s) correlation. The width of the cluster component is = 2.2680 Å. The correlation (g t) is shown in the lower panel of the same figure. In this case the sky emission affects only the cluster spectrum and hence the sky component does not show up. The cluster component is = 2.1147 Å.
@@
We have combined these results, together with the width of the (t t) correlation function ( = 1.241 Å) and we have applied the three methods studied in Zamora & Díaz (2023c):
@@
the last equation being the one proposed by Tonry & Davis (1979). We have performed theses calculations for R52 correcting methods (i) and (ii) for the non-Gaussian behaviour of the (t t) correlation. The final results are (i) = 27.87 km/s; (ii) = 28.42 km/s; and (iii) = 28.44 km/s. We can see that the three methods show very similar results, thus we can conclude that the reduction and treatment of these high resolution data and the application of the cross-correlation method is correct.
@@
Therefore, for the rest of CNSFRs, we have decided to apply method (iii) in order not to use the (g g) correlation due to the sky subtraction problems. The velocity dispersion errors have been calculated as the semi-difference between the largest and smallest Gaussian width that can be fitted considering the asymmetries in the correlation peak as suggested in Zamora & Díaz (2023c), although it could lead to the error being somewhat overestimated. Applying the described methodology, we have obtained reliable CaT velocity dispersion values for 30 of the selected CNSFR regions. They are shown in Table 6 and they take values from 14.3 to 77.1 km/s.
@@ 4.1 Cluster sizes
We have calculated the effective radii of the star clusters from the Hubble Space Telescope (HST) continuum image. These data were acquired on 2009 November 5 with the Wide Field and Planetary Camera 2 (WFPC2) and the F555W continuum filter where stellar clusters are easily identified. These data are a part of proposal 6276 (IP: J. Westphal) organised in 3 exposures of 480 s. The data have been downloaded from the Hubble Legacy Archive and reduced by the Space Telescope Science Institute (STScI) using available calibration files taken for this observation and keeping in mind different dithering positions.
@@
The effective radii of the star clusters have been calculated using the following procedure. First, we have estimated the region background by fitting a three-order polynomial. After subtracting it, we fitted the surface brightness of each knot present in the observed clusters assuming a two-dimensional Gaussian profile. Fig. 8 shows two examples of the described procedure as applied to regions R1, and R52. The left panels of the figure show the F555W WFPC2-HST image for each selected region and the right panels show the same image after background subtraction. Selected clusters are shown by blue circles. The radius of each knot has been taken as 1/2 FWHM. There are three regions, R72, R75, and R88, where no knot is readily appreciable in the continuum at 5550 Å thus, we have not calculated their sizes.
@@
As opposed to other CNSFRs from the literature, this galaxy does not seem to have large star-forming complexes and 60 % of the CNSFRs (16 out of 27) seem to host single clusters and 27 % (7 out of 27) are composed by two. Three regions show three different knots, and one shows four. We can compare these values with the results obtained for NGC 7469 (see Zamora & Díaz, 2023a) which shows two regions with 20 and 19 knots respectively and only 15 % of the regions (4 of 27) seem to have only one star cluster. These later results, are similar to those found for the regions analysed in NGC 2903 showing more than 20 knots (Hägele et al., 2009), NGC 3310 hosting two great complexes with 28 and 30 clusters each (Hägele et al., 2010) and NGC 3351 with a large region showing 17 distinct knots (Hägele et al., 2007). In the case of NGC 7469 one can imagine this effect to be due to the galaxy being further away since in this case we may be selecting larger star-forming complexes due to lack of spatial resolution; however, the distances to the other three galaxies mentioned are of the same order as the one studied here. In particular, NGC 3310 at a similar distance to NGC 7742 seems to have only one region hosting an individual knot. It is probable that the difference between the CNSFR is related to their different formation processes which in the latter case seem to be due to gas accretion after a minor merger event.
@@
The effective radii of the studied single knots range between 8.5 and 60.5 pc, with a mean value of 12.6 pc. Figure 9 shows the distribution of the radius of each knot for the CNSFRs in this work and those found in the literature used for comparison (Hägele et al., 2009, 2007, 2010). The range of radius found for NGC 7469 is wider than in the other galaxies and the same is true for the mean radii for the inner and outer ring regions. On the other hand, NGC 3351, NGC 2903, and NGC 3310 have smaller knots than NGC 7742.

### v3

@@ 3.3 Stellar absorption lines
The main objective of our work is to estimate the properties of the stellar population of the clusters associated with the observed CNSFRs. In order to do this, we used two sets of absorption lines: the MgIb triplet at 5167-5183 detected in the MUSE data (unresolved) and the CaII triplet (CaT) at 8498, 8542, 8662 Å (CaT) resolved with MEGARA.
We measured the equivalent widths (EWs) of these stellar absorption lines by integrating the line fluxes and taking as continuum values the linear interpolation between the mean integrated fluxes in 30 Å continuum bands on both sides of the line. The details of the measuring procedure can be seen in Table 8 in Zamora & Díaz (2023a). The EW errors were calculated using the standard deviation in the two continuum bands and propagating them in quadrature. In principle, absorption line EW measurements in objects with different velocity dispersions should be for the broadening of the spectral lines; this effect decreases the continuum level providing lower EW values. However, this correction for MUSE data was found to be negligible (¡ 1 Å for velocity dispersions of around 170 km/s for MUSE spectral resolution), and so no correction was applied. Table 5 lists in columns 1 to 3: (1) the region ID; (2) the CaT EW in Å; and (3) the MgIb EW also in Å .
@@
The presence of a non-negligible contribution by the nebular continuum can dilute the starlight weakening and distorting the absorption features. This component is usually not taken into account, but it provides information about the origin of the features measured. Thus, we calculated a dilution factor, D, for each set of lines with respect to a reference average value measured in the spectra of normal spiral galaxy nuclei, which corresponds to old metal-rich stellar populations (EWref(CaII) = 7.7 0.5 Å, EWref(MgI) = 5.18 0.71 Å; Terlevich et al., 1990; Keel, 1983). We calculated this dilution factor as the ratio of the measured EWs to the adopted reference values, D = EWobs / EWref, and compared it to the dilution produced by the nebular continuum associated with a nebula ionised by a young star cluster synthesised using the PopStar code (Mollá et al., 2009) with Salpeter’s IMF (mlow = 0.85 M⊙, mup = 120 M⊙ Salpeter, 1955) and an age of 5.5 Ma.
Figure 4 shows the comparison between the dilution factors, D, calculated for the MgIb and CaT stellar absorption features for the studied CNSFRs. The red lines represent the dilution produced by the nebular continuum in the optical and near-IR. We can see that the mean value of the MgIb dilution is about 50%, consistent with the contribution by a nebular continuum. However, the CaT lines are shown almost undiluted. This result can be explained by the presence of red supergiant stars whose CaT features are stronger than those in normal early spiral galaxies, while the reference value probably consists of a stellar population dominated by red giants. This points to the occurrence of a relatively recent star formation episode. The same behaviour has already been found for circumnuclear star clusters in both rings of NGC 7469 (see Zamora & Díaz, 2023a) and in earlier work (Terlevich et al., 1990; Garcia Vargas et al., 1993; Oliva et al., 1995).
@@ 3.4 Stellar velocity dispersions
Stellar velocity dispersions can be obtained from the stellar absorption lines. In the second paper of this series (see Zamora & Díaz, 2023a) we used the CaT lines detected in the MUSE data for this calculation. However, the spectral resolution reached in the MUSE data was insufficient to resolve the CaT lines in the case of the HII ring regions in NGC 7742, and therefore we acquired higher resolution data using MEGARA, which provides about 14 km/s of spectral resolution (see Sect. 2). As in the case of NGC 7469, the clusters studied in this work have red supergiant stars dominating the near-IR range of the spectrum and the CaT 8498, 8542, 8662 Å lines (see Sect. 3.3) are easily detected. The use of the near-IR spectra provides two principal advantages for this study: there is little contamination by TiO bands and nebular lines, and the velocity resolution at longer wavelengths is higher than in the blue part of the spectrum for the same spectral dispersion. Additionally, using the cross-correlation technique, stellar weak lines present in this part of the spectra also contribute to the goodness of the results.
@@
We used the cross-correlation technique proposed by Tonry & Davis (1979). In this traditional method, a comparison between a stellar template with the observed stellar population spectrum is done. The assumption made assumes that a galaxy spectrum can be represented by the sum of different stellar spectra with different velocity offsets convolved with a broadening function: , where b is the broadening function; g and t are the cluster and the stellar template spectra, respectively; and is the velocity shift between the stars. The width of this broadening function is calculated using two correlations: the cluster spectrum with the stellar template (g t) and the stellar template with itself (t t). The convolution products are applied assuming a periodic spectrum with discrete Fourier transforms, and the broadening function is assumed to be Gaussian. However, this last assumption is not correct for the (t t) correlation at spectral resolutions as high as the one provided by MEGARA, and this technique needs to be modified. We used the adapted methods proposed in Zamora & Díaz (2023c) of the (g g) correlation, which is broader than the other two and whose Gaussian behaviour can be ensured. In order to test these two additional methods and the traditional ones with our data, we selected a region with high S/N (R52, located in the pointing with the higher exposure time in the HR-I set-up). For the stellar template, we used late-type red giants and supergiants obtained from the MEGARA stellar library (García-Vargas et al., 2020) aligning all available stellar spectra in velocity and calculating the average between them, verifying that no apparent broadening is introduced in the procedure. The use of only two stellar types of stars for the correlation analysis could introduce errors in our velocity dispersion measures. However, the use of CaT lines minimises the mismatch between the stellar template and the cluster spectrum since they are very strong in most stars of moderate to high metallicity.
@@
The top panel of Fig. 5 shows the stellar spectrum of the selected cluster R52 at high and low MEGARA spectral resolutions, with red and purple lines respectively. The stellar template is shown with a blue line and the nebular lines present in this spectral range are marked. We can see that the S/N values of these spectra are low, even in this region that was observed with twice the average exposure time. This could be due to the noise component or to errors associated with the subtraction of the sky background, which is very prominent in the near-IR. Thus, we calculated a mean spectrum of the sky, using the MEGARA fibres earmarked for this purpose. The lower panel of the figure shows this spectrum, which shares some features with the cluster spectrum.
@@
We used the modified version of the Tonry & Davis method described above for the application of the cross-correlation technique to the data obtained from the MEGARA in the HR-I configuration in the spectral range 8450 Åto 8850 Å. Prior to this, we binned the spectrum into 4069 logarithmic wavelength bins corresponding to a velocity resolution of 3.4 km/s. Then, we estimated the continuum by fitting a second-order polynomial after masking the features present in the spectrum. The masks were built assuming a width of 3 Å and 10 Å on each side of the central wavelength of the nebular and stellar lines involved, respectively. This continuum was subsequently subtracted.
@@
Next, the high- and low-frequency variations were filtered using a band-pass filter in the Fourier spectrum transform. The minimum and maximum wave numbers are kmin = 3 and kmax = 350, which correspond to wavelength values lower than 10 Å and to the nominal MEGARA/HR-I spectral resolution. The low-frequency variations are associated with continuum subtracted errors and the high-frequency variations with the noise component. Figure 6 shows, in a small range of wavelengths, in light red the original spectrum and in dark red the spectrum after the band-pass filter. The stellar template is also shown in blue for comparison. We can see that this filter removes part of the noise, recovering some stellar features. However, the S/N continues to be low and we can see residuals, for example in the central part of the strongest line. Thus, the noise present in our spectrum probably comes from the errors associated with the pipeline sky subtraction. Therefore, we decided to use the data without the sky subtraction to be able to quantify this additional component in the cross-correlation technique and to obtain a more reliable result in the stellar dispersion determination.
@@
An important step in the cross-correlation procedure is the subtraction of the non-stellar component of the spectra, which can introduce errors in the velocity dispersions of up to 5 km/s (see Zamora & Díaz, 2023c). In order to remove correctly the nebular component of the spectrum, we first removed the proper motion of the cluster by shifting the spectrum to the rest-frame position. For this step, we used the position of the (g t) cross-correlation. Then, we masked the nebular lines shown in Fig. 5 assuming a width of 0.4 Å at each side of their central wavelength.
@@
Finally, we calculated four cross-correlation functions for this HII region, (t t), (g t), (g g), and (s s), with t being the stellar template, g the cluster spectrum, and s the mean sky spectrum of our observations. The top panel of Fig. 7 shows the (s s) correlation. Two components can be seen, one associated with the sky lines and another associated with the detector response. They were fitted with two Gaussian functions and their widths are = 0.2098 Å and = 12.1445 Å, respectively. The width of the first component corresponds to a value of = 0.15 Å in -units, which is slightly lower than the MEGARA resolution for this set-up, = 0.18.
@@
The top panel of Fig. 7 shows, as an example, the (g g) correlation for the R52 cluster. In this figure we can identify the component associated with the sky lines and the component associated with the stars of the cluster. According to this, we fitted simultaneously two Gaussian functions, one of them using the value calculated for the (s s) correlation. The width of the cluster component is = 2.2680 Å. The correlation (g t) is shown in the lower panel of the same figure. In this case the sky emission affects only the cluster spectrum, and hence the sky component does not show up. The cluster component is = 2.1147 Å.
@@
We combined these results, together with the width of the (t t) correlation function ( = 1.241 Å) and we applied the three methods studied in Zamora & Díaz (2023c):
@@
The last equation is the one proposed by Tonry & Davis (1979). We performed these calculations for R52, correcting methods (i) and (ii) for the non-Gaussian behaviour of the (t t) correlation. The final results are (i) = 27.87 km/s; (ii) = 28.42 km/s; and (iii) = 28.44 km/s. We can see that the three methods show very similar results, and thus we can conclude that the reduction and treatment of these high-resolution data and the application of the cross-correlation method is correct.
@@
Therefore, for the rest of the CNSFRs, we decided to apply method (iii) in order not to use the (g g) correlation, due to the sky subtraction problems. The velocity dispersion errors were calculated as the semi-difference between the largest and smallest Gaussian width that can be fitted considering the asymmetries in the correlation peak, as suggested in Zamora & Díaz (2023c), although it could lead to the error being somewhat overestimated. Applying the described methodology, we obtained reliable CaT velocity dispersion values for 30 of the selected CNSFR regions. They are shown in Table 6, and they take values from 14.3 to 77.1 km/s.
@@ 4.1 Cluster sizes
We calculated the effective radii of the star clusters from the Hubble Space Telescope (HST) continuum image. These data were acquired on 2009 November 5 with the Wide Field and Planetary Camera 2 (WFPC2) and the F555W continuum filter where stellar clusters are easily identified. These data are a part of proposal 6276 (IP: J. Westphal) organised in three exposures of 480 s. The data were downloaded from the Hubble Legacy Archive and reduced by the Space Telescope Science Institute (STScI) using the available calibration files taken for this observation and keeping in mind different dithering positions.
@@
The effective radii of the star clusters were calculated using the following procedure. First, we estimated the region background by fitting a three-order polynomial. After subtracting it, we fitted the surface brightness of each knot present in the observed clusters assuming a two-dimensional Gaussian profile. Figure 8 shows two examples of the described procedure as applied to regions R1, and R52. The left panels of the figure show the F555W WFPC2-HST image for each selected region and the right panels show the same image after background subtraction. Selected clusters are shown by blue circles. The radius of each knot is taken as 1/2 FWHM. There are three regions, R72, R75, and R88, where no knot is readily appreciable in the continuum at 5550 Å, and thus we did not calculate their sizes.
@@
As opposed to other CNSFRs from the literature, this galaxy does not seem to have large star-forming complexes, and 60 % of the CNSFRs (16 out of 27) seem to host single clusters and 27 % (7 out of 27) are composed of two clusters. Three regions show three different knots, and one shows four. We can compare these values with the results obtained for NGC 7469 (see Zamora & Díaz, 2023a), which shows two regions with respectively 20 and 19 knots, and only 15 % of the regions (4 of 27) seem to have only one star cluster. These last results, are similar to those found for the regions analysed in NGC 2903 showing more than 20 knots (Hägele et al., 2009), NGC 3310 hosting two great complexes with 28 and 30 clusters each (Hägele et al., 2010), and NGC 3351 with a large region showing 17 distinct knots (Hägele et al., 2007). In the case of NGC 7469, we imagine this effect to be due to the galaxy being further away since in this case we may be selecting larger star-forming complexes, due to lack of spatial resolution; however, the distances to the other three galaxies mentioned are of the same order as the one studied here. In particular, NGC 3310 at a similar distance to NGC 7742 seems to have only one region hosting an individual knot. It is probable that the difference between the CNSFRs is related to their different formation processes which in the latter case seem to be due to gas accretion after a minor merger event.
@@
The effective radii of the studied single knots range between 8.5 and 60.5 pc, with a mean value of 12.6 pc. Figure 9 shows the distribution of the radius of each knot for the CNSFRs in this work and those found in the literature used for comparison (Hägele et al., 2009, 2007, 2010). The range of radius found for NGC 7469 is wider than in the other galaxies, and the same is true for the mean radii for the inner and outer ring regions. On the other hand, NGC 3351, NGC 2903, and NGC 3310 have smaller knots than NGC 7742.